Janus Products, Inc. is a merchandising company that sells binders, paper, and other school supplies. The company is planningRequired: 1. Prepare a schedule of expected cash collections for July, August, and September and for the quarter in total. JAb. A schedule of expected cash disbursements for merchandise purchases for July, August, and September and for the quarter in3. Prepare a cash budget for July, August, and September and for the quarter in total. (Roundup Borrowing and Repayments
Answer #1

1.

Schedule of Expected Cash Collections
July August September Quarter
Cash Sales (20%) $          9,400 $     15,400 $         11,400 $        36,200
Credit Sales
May $          5,920 $           5,920
June $        24,080 $       6,880 $        30,960
July $          3,760 $     26,320 $           7,520 $        37,600
August $       6,160 $         43,120 $        49,280
September $           4,560 $           4,560
Total Cash Collections $        43,160 $     54,760 $         66,600 $      164,520

Working

July August September
Cash Sales (20%) =47000*20% =77000*20% =57000*20%
Credit Sales
May =37000*80%*20%
June =43000*80%*70% =43000*80%*20%
July =47000*80%*10% =47000*80%*70% =47000*80%*20%
August =77000*80%*10% =77000*80%*70%
September =57000*80%*10%

2.a.

Merchandise Purchase Budget
July August September
Cost of Goods Sold $        26,800 $     44,800 $         32,800
Add : Desired Ending Inventory $        33,600 $     24,600 $         22,350
Total Needs $        60,400 $     69,400 $         55,150
Less Beginning Inventory $        21,500 $     33,600 $         24,600
Merchandise Purchase $        38,900 $     35,800 $         30,550


Ending Inventory is 75% of next month sales
Beginning Inventory is ending inventory of previous month

2b.

July August September Quarter
Accounts Payable, June 30 $        15,200
July Purchases $        19,450 $     19,450
August Purchases $     17,900 $         17,900
September Purchases $         15,275
Total Cash Disbursements $        34,650 $     37,350 $         33,175 $      105,175

3.

Cash Budget
July August September Quarter
Beginning Cash $          8,700 $       9,410 $           9,170 $           8,700
Total Cash Collections $        43,160 $     54,760 $         66,600 $      164,520
Total Cash Available $        51,860 $     64,170 $         75,770 $      173,220
Deduct : Disbursements
For Merchandise purchase $        34,650 $     37,350 $         33,175 $      105,175
Selling Expense $          9,300 $     13,100 $           9,200 $        31,600
Administrative expense $          3,650 $       5,550 $           4,450 $        13,650
Purchase of Land $          4,850 $           4,850
Dividend Paid $           1,350 $           1,350
Total Disbursements $        52,450 $     56,000 $         48,175 $      156,625
Excess (Deficiency) of cash $            -590 $       8,170 $         27,595 $        16,595
Financing
Borrowing $        10,000 $       1,000 $        11,000
Interest $             -320 $             -320
Repayment $       -11,000 $       -11,000
Total Financing $        10,000 $       1,000 $       -11,320 $             -320
Ending Balance $          9,410 $       9,170 $         16,275 $        16,275

Cash Paid for Administrative expense = Administrative expense - Depreciation expense
Interest paid = $10000 x 3% + $1000 x 2% =$320